### v1

@@ 4.1 Strategy of the proof
$H=H(n,\delta):=\left\lfloor\frac{h(q_{n})}{\lfloor\delta n\rfloor}\right\rfloor.$ (58)
The interval $\big[s\lfloor\delta n\rfloor,(s+1)\lfloor\delta n\rfloor\big)$ is called the $s$-th height interval. We focus on the coalescent process starting from height $(s+1)\lfloor\delta n\rfloor$ for $s=1,\dots,H-2$. Recall that $\mathcal{X}_{s\lfloor\delta n\rfloor}\big(n,(s+1)\lfloor\delta n\rfloor\big)$ consists of all vertices at height $s\lfloor\delta n\rfloor$ that have at least one descendant at height $(s+1)\lfloor\delta n\rfloor$. Those vertices are referred to as the $\delta$-coalescent vertices at the $s$-th height interval. The collection of all $\delta$-coalescent vertices over these height intervals is denoted by
@@
These $\delta$-coalescent vertices serve as structural anchors within the tree, ensuring that every vertex lies within a distance of at most $3\lfloor\delta n\rfloor$ from one of these vertices. (In general, the distance bound is $\lfloor\delta n\rfloor$, but it increases to $2\lfloor\delta n\rfloor$ near the root since $s=0$ is not considered, and up to $3\lfloor\delta n\rfloor$ near height $h(q_{n})$ as s=K1s=K-1 is also excluded.)
@@
    each $\delta$-coalescent vertex has at least $\Theta(n^{2})$ descendants within distance $\lfloor\delta n\rfloor$ (see Proposition 4.5).
These two results imply that, as $k\to\infty$, with high probability, every $\delta$-coalescent vertex will have a descendant among the $k$ selected vertices $\{V_{n,1},\ldots,V_{n,k}\}$ within its $\lfloor\delta n\rfloor$-neighborhood. Consequently, $\frac{1}{n}\Delta(n,k)$ remains small as $n\to\infty$ and $k\to\infty$.
We now formulate the two results as propositions.
@@
Assume (H1) and (H2’). Then for any $\delta>0$, the number of $\delta$-coalescent vertices, $\bar{\mathcal{X}}(n,\delta)$, is tight, i.e.,
@@
This follows directly from Theorem 3.26, with $h^{*}=(s+1)\lfloor\delta n\rfloor$ and $h_{*}=s\lfloor\delta n\rfloor$.
@@
$B(n,\gamma,\delta):=\Big\{\forall\,v\in\bar{\mathcal{X}}(n,\delta):\ \#\{u\in\text{Desc}(v):d(u,v)<\delta n\}\geq\gamma n^{2}\Big\}.$ (59)
@@
$A(n,k,\delta):=\big\{\forall\,v\in\bar{\mathcal{X}}(n,\delta),\ \exists\,1\leq i\leq k\text{ such that }V_{n,i}\in\text{Desc}(v)\text{ and }d(v,V_{n,i})\leq\delta n\big\}.$
By (2), there exists constant $C>0$ such that
$V(\mathscr{T}^{n})\leq Cn^{2}.$
It follows that for any $M\geq 1$ and $\gamma>0$,
@@
Applying Propositions 4.3 and 4.5, we conclude that
@@
$A(n,k,\delta)\subset\big\{\Delta(n,k)\leq 4\delta n\big\}.$ (60)
Indeed, since the root $v_{0}$ is a vertex in $\mathscr{T}^{n}_{k}$, for any vertex $v$ with $\left\lvert v\right\rvert\in\big[0,4\delta n\big]$, we have
$\min_{u\in V(\mathscr{T}^{n}_{k})}d(v,u)\leq d(v,v_{0})\leq 4\delta n.$
For all vertex $v$ with $\left\lvert v\right\rvert\in\big(s\delta n,(s+1)\delta n\big]$ for some $s=4,\dots,H$, let $v^{*}$ be its ancestor at height $(s-2)\lfloor\delta n\rfloor$. By definition, $v^{*}\in\bar{\mathcal{X}}(n,\delta)$. Moreover, on the event $A(n,k,\delta)$, there exists $1\leq i_{0}\leq k$ such that $V_{n,i_{0}}\in\text{Desc}(v^{*})$ and $d(v^{*},V_{n,i_{0}})<\delta n$. Consequently,
$\min_{u\in V(\mathscr{T}^{n}_{k})}d(v,u)\leq d(v,V_{n,i_{0}})\leq d(v,v^{*})+d(v^{*},V_{n,i_{0}})\leq 4\delta n,\quad\text{on }A(n,k,\delta).$
This verifies (60) and thus concludes the proof.
@@
Recall the definition of $B(n,\eta,\delta)$ from (59). For $s=1,\ldots,H-2$, define the event
$B(n,s,\gamma,\delta):=\Big\{\forall\,v\in X_{s\lfloor\delta n\rfloor}\big(n,(s+1)\lfloor\delta n\rfloor\big),\ \#\{u\in\text{Desc}(v):d(u,v)<\delta n\}\geq\gamma n^{2}\Big\}.$ (61)
This represents the “restriction” of $B(n,\gamma,\delta)$ to the $s$-th height interval. It suffices to fix $\delta>0$ and $1\leq s\leq H-2$, and show that for any $\eta>0$, there exists $\gamma=\gamma(\eta)>0$ such that for all $n\geq 1$,
${\mathbb{P}}_{n}\big(B(n,s,\gamma,\delta)\big)\geq 1-\eta.$ (62)
Our strategy is as follows. Recall the base height $h^{*}$ in the definitions of $\mathcal{X}_{j}=\mathcal{X}_{j}(n,h^{*})$ and $X_{j}=X_{j}(n,h^{*})$. In this subsection we fix
@@
and consider the coalescent process $(X_{j})_{0\leq j\leq\lfloor\delta n\rfloor}$ from this base height. We first analyze the behavior of the coalescent process $(X_{j})$ and identify a typical event that occurs with high probability and is more tractable for subsequent estimates. We then condition on this event and establish the desired bound for the probability of $B(n,s,\gamma,\delta)$.
@@
${\mathbb{P}}_{n}\!\Big(X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}\leq M_{0}\Big)\geq 1-\eta/4,\quad\text{for all }n\geq 1.$ (63)
Next, consider the holding time
$\bar{T}=\bar{T}(n,\delta):=\inf\Big\{j\geq 1:X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor-j}<X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}\Big\}.$
Using the same argument as in (46) together with (34), we obtain:
@@
There exists a constant $C=C(\delta,\varepsilon,M_{0})>0$ such that for any $1\leq m\leq M_{0}$, conditionally on $X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}=m$, we have
$\bar{T}\;\geq_{\rm sto}\;\text{Geometric}(C/n)\wedge\tfrac{1}{4}\delta n.$
Consequently, there exists $c_{0}=c_{0}(C,\delta)\in(0,\delta/2)$ such that for any $1\leq m\leq M_{0}$,
${\mathbb{P}}_{n}\!\big(\bar{T}>\lfloor c_{0}n\rfloor\,\big|\,X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}=m\big)\geq 1-\eta/4,\quad\text{for all }n\geq 1.$ (64)
@@
$\bar{h}:=(s+1)\lfloor\delta n\rfloor-\left\lfloor\frac{1}{2}\delta n\right\rfloor-\lfloor c_{0}n\rfloor.$
Combining (63) and (64), we arrive at the following high-probability event:
$\begin{aligned} &E(n,\delta,\eta,m):=\Big\{\bar{T}>\lfloor c_{0}n\rfloor,X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}=m\Big\}=\Big\{X_{k}=m,\ \forall\,k\in[\,\bar{h},\bar{h}+\lfloor c_{0}n\rfloor]\Big\},\\ &{\mathbb{P}}_{n}\bigg(\bigcup_{m=1}^{M_{0}}E(n,\delta,\eta,m)\bigg)\geq 1-\eta/2.\end{aligned}$ (65)
We shall take the event $E(n,\delta,\eta,m)$ as the typical event on which our subsequent analysis will be conditioned. Define for $1\leq m\leq M_{0}$ the conditional probability measure
${\mathbb{Q}}^{m}_{n}(\cdot):={\mathbb{P}}_{n}\big(\,\cdot\,\big|\,E(n,\delta,\eta,m)\big).$
Structure of $\mathscr{T}^{n}$ on $E(n,\delta,\eta,m)$
The purpose of this subsection is to set up the structural ingredients — coalescent subtrees, trunk paths and vertices — which will be crucial for the forthcoming estimates. Recall the subtree notation $\mathscr{T}^{n}(v)=\mathscr{T}^{n}(v,h^{*})$ from (26), which denotes the subtree of $\mathscr{T}^{n}$ rooted at $v$ and truncated at height $h^{*}$.
@@
$\Lambda_{j}:\mathbf{T}_{j}\longrightarrow\{1,\ldots,|\mathbf{T}_{j}|\},\qquad\Lambda_{j}(\mathsf{T}_{i}):=i,$ (66)
@@
In other words, for each $j$ between $\bar{h}$ and $\bar{h}+\lfloor c_{0}n\rfloor$, there are exactly $m$ vertices in $\mathcal{X}_{j}$. These vertices belong to $m$ distinct ancestral lineages that remain separate and do not coalesce.
Let $\mathsf{T}_{1},\ldots,\mathsf{T}_{m}$ be the $m$ coalescent subtrees at height $\bar{h}$. For $\bar{h}\leq j\leq\bar{h}+\lfloor c_{0}n\rfloor$ and $1\leq i\leq m$, denote by
$z^{j}_{i}\in V(\mathsf{T}_{i})$
the unique vertex of $\mathsf{T}_{i}$ at height $j$. The sequence
@@
is called the trunk path of $\mathsf{T}_{i}$, and its elements are referred to as the trunk vertices of $\mathsf{T}_{i}$.
The remaining $q_{n}(j)-m$ non-trunk vertices at height $j$ are denoted by $y^{j}_{1},\ldots,y^{j}_{q_{n}(j)-m}$, where the subscripts are assigned arbitrary.
@@
Assume (H1) and (H2’). For fixed $\delta,s,\eta,M_{0},c_{0}$, there exists $\gamma>0$ such that for all $1\leq m\leq M_{0}$ and sufficiently large $n$,
${\mathbb{Q}}^{m}_{n}\big(\mathcal{N}_{i}\geq\gamma n^{2},\ \forall 1\leq i\leq m\big)\geq 1-\eta/2.$ (67)
Proof 4.10 (Proof of (62) assuming Theorem 4.9).
Each $\delta$-coalescent vertex at height $s\lfloor\delta n\rfloor$ has at least one descendant in $\{z^{0}_{1},\dots,z^{0}_{m}\}$. Hence, for all large $n$ and $1\leq m\leq M_{0}$,
${\mathbb{Q}}^{m}_{n}\big(B(n,s,\gamma,\delta)\big)\geq{\mathbb{Q}}^{m}_{n}\big(\mathcal{N}_{i}\geq\gamma n^{2},\ \forall 1\leq i\leq m\big)\geq 1-\eta/2.$
Combining with (65) completes the proof of (62).
@@ 4.4.1 Outline of the proof
The first step is to partition the interval $\big[\bar{h},\bar{h}+\lfloor c_{0}n\rfloor\big]$ into $K_{\theta}$ sub-intervals $(h_{r},h_{r-1}]$, where $\bar{h}+\lfloor c_{0}n\rfloor=h_{0}>h_{1}>\cdots>h_{K_{\theta}}$ and each sub-interval has length $\lfloor\theta n\rfloor$ (see (68) below for precise definitions). For each $r=0,\ldots,K_{\theta}-1$, let $\mathcal{N}_{i,r}$ be the number of $y^{j}_{l}$ in the height interval $(h_{r},h_{r-1}]$ that merge into $Z_{i}$ before height $h_{r}$.
We then establish the independence of the family $(\mathcal{N}_{i,r})$ under $\mathbb{Q}^{m}_{n}$ (Proposition 4.11) and obtain a uniform lower bound for their tail probabilities (Proposition 4.12). Since $\mathcal{N}_{i}$ exceeds the sum of $\mathcal{N}_{i,r}$, the event that $\mathcal{N}_{i}>\gamma n^{2}$ is implied by the event that some $\mathcal{N}_{i,r}>\gamma n^{2}$. The independence and uniform lower bound together yield the desired probability estimate in (67).
@@
For these events, we prove a general lemma (Lemma 4.15) showing that their conditional probabilities can be decomposed into contributions from the genealogies above and below a given height. This decomposition directly yields the desired independence property and simplifies the probability estimates required for proving the uniform lower bound.

### v2

@@ 4.1 Strategy of the proof
$H=H(n,\delta):=\left\lfloor\frac{h(q_{n})}{\lfloor\delta n\rfloor}\right\rfloor.$ (4.1)
The interval $\big[s\lfloor\delta n\rfloor,(s+1)\lfloor\delta n\rfloor\big)$ is called the $s$-th height interval. We focus on the coalescent process starting from height $(s+1)\lfloor\delta n\rfloor$ for $s=1,\dots,H-2$. Recall that $\mathcal{X}_{s\lfloor\delta n\rfloor}\big(n,(s+1)\lfloor\delta n\rfloor\big)$ consists of all vertices at height $s\lfloor\delta n\rfloor$ that have at least one descendant at height $(s+1)\lfloor\delta n\rfloor$. Those vertices are referred to as the $\delta$-coalescent vertices at the $s$-th height interval. The collection of all $\delta$-coalescent vertices at these height intervals is denoted by
@@
These $\delta$-coalescent vertices serve as structural anchors within the tree, ensuring that every vertex lies within a distance of at most $3\lfloor\delta n\rfloor$ from one of these vertices. (In general, the distance bound is $\lfloor\delta n\rfloor$, but it increases to $2\lfloor\delta n\rfloor$ near the root since $s=0$ is not considered, and up to $3\lfloor\delta n\rfloor$ near height $h(q_{n})$ as s=H1s=H-1 is also excluded.)
@@
    with probability arbitrarily close to 1, each $\delta$-coalescent vertex has at least $\Theta(n^{2})$ descendants within distance $\lfloor\delta n\rfloor$ (see Proposition 4.4).
These two results imply that, for all $k$ sufficiently large, with high probability, every $\delta$-coalescent vertex will have a descendant among the $k$ selected vertices $\{V^{n}_{1},\ldots,V^{n}_{k}\}$ within its $\lfloor\delta n\rfloor$-neighborhood. Consequently, $\frac{1}{n}\Delta(n,k)$ remains small as $n\to\infty$ and $k\to\infty$.
We now formulate the two results as the following propositions.
@@
Assume (H1) and (H2a). Then for any $\delta>0$, the number of $\delta$-coalescent vertices, $\bar{\mathcal{X}}(n,\delta)$, is tight, i.e.,
@@
This follows directly from Theorem 3.13, with $h^{*}=(s+1)\lfloor\delta n\rfloor$ and $h_{*}=s\lfloor\delta n\rfloor$. ∎
@@
$B(n,\gamma,\delta):=\Big\{\forall\,v\in\bar{\mathcal{X}}(n,\delta):\ \#\{u\in\text{Desc}(v):d(u,v)<\delta n\}\geq\gamma n^{2}\Big\}.$ (4.2)
@@
$A(n,k,\delta):=\big\{\forall\,v\in\bar{\mathcal{X}}(n,\delta),\ \exists\,1\leq i\leq k\text{ such that }V^{n}_{i}\in\text{Desc}(v)\text{ and }d(v,V^{n}_{i})\leq\delta n\big\}.$
By (1.1), there exists a constant $C>0$ such that $V(\mathscr{T}^{n})\leq Cn^{2}.$ It follows that for any $M\geq 1$ and $\gamma>0$,
@@
Combining Propositions 4.3 and 4.4, we conclude that
@@
$A(n,k,\delta)\subset\big\{\Delta(n,k)\leq 4\delta n\big\}.$ (4.3)
Indeed, since the root $\rho$ is a vertex in $\mathscr{T}^{n}_{k}$, for any vertex $v$ with $\left\lvert v\right\rvert\in\big[0,4\delta n\big]$, we have
$\min_{u\in V(\mathscr{T}^{n}_{k})}d(v,u)\leq d(v,\rho)\leq 4\delta n.$
For all vertex $v$ with $\left\lvert v\right\rvert\in\big(s\delta n,(s+1)\delta n\big]$ for some $s=4,\dots,H$, let $v^{*}$ be its ancestor at height $(s-2)\lfloor\delta n\rfloor$. By definition, $v^{*}\in\bar{\mathcal{X}}(n,\delta)$. Moreover, on the event $A(n,k,\delta)$, there exists $1\leq i_{0}\leq k$ such that $V^{n}_{i_{0}}\in\text{Desc}(v^{*})$ and $d(v^{*},V^{n}_{i_{0}})<\delta n$. Consequently,
$\min_{u\in V(\mathscr{T}^{n}_{k})}d(v,u)\leq d(v,V^{n}_{i_{0}})\leq d(v,v^{*})+d(v^{*},V^{n}_{i_{0}})\leq 4\delta n,\quad\text{on }A(n,k,\delta).$
This verifies (4.3) and thus concludes the proof.
@@
Recall the definition of $B(n,\gamma,\delta)$ from (4.2). For $s=1,\ldots,H-2$, define the event
$B(n,s,\gamma,\delta):=\Big\{\forall\,v\in X_{s\lfloor\delta n\rfloor}\big(n,(s+1)\lfloor\delta n\rfloor\big),\ \#\{u\in\text{Desc}(v):d(u,v)<\delta n\}\geq\gamma n^{2}\Big\}.$ (4.4)
This represents the “restriction” of $B(n,\gamma,\delta)$ to the $s$-th height interval. It suffices to fix $\delta>0$ and $1\leq s\leq H-2$, and show that for any $\eta>0$, there exists $\gamma=\gamma(\eta)>0$ such that for all $n$ sufficiently large,
${\mathbb{P}}_{n}\big(B(n,s,\gamma,\delta)\big)\geq 1-\eta.$ (4.5)
Our strategy is as follows. Recall the base height $h^{*}$ in the definitions of $\mathcal{X}_{j}=\mathcal{X}_{j}(n,h^{*})$ and $X_{j}=X_{j}(n,h^{*})$. In this subsection, we fix
@@
and consider the coalescent process $(X_{j}:h^{*}-\lfloor\delta n\rfloor\leq j\leq h^{*})$ from this base height. We first analyze the behavior of the coalescent process $(X_{j})$ and identify a typical event that occurs with high probability and is more tractable for subsequent estimates. We then condition on this event and establish the desired bound for the probability of $B(n,s,\gamma,\delta)$.
@@
${\mathbb{P}}_{n}\!\Big(X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}\leq M_{0}\Big)\geq 1-\tfrac{\eta}{4},\quad\text{for all }n\geq 1.$ (4.6)
Next, consider
$T=T(n,\delta):=\sup\Big\{0\leq j<h^{*}-\lfloor\tfrac{1}{2}\delta n\rfloor:X_{j}<X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}\Big\}.$
Using the same argument as in (3.39) together with (3.26), we obtain:
@@
There exists a constant $C=C(\delta,\varepsilon,M_{0})>0$ such that for any $1\leq m\leq M_{0}$, conditionally on $X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}=m$,
$h^{*}-\lfloor\tfrac{1}{2}\delta n\rfloor-T\;\geq_{\rm sto}\;{\rm Geometric}\bigg(\frac{C}{n}\bigg)\wedge\frac{1}{4}\delta n.$
Consequently, for any $\eta>0$, there exists $c_{0}=c_{0}(C,\delta,\eta)\in(0,\delta/2)$ such that for all $1\leq m\leq M_{0}$,
${\mathbb{P}}_{n}\!\Big(T\leq h^{*}-\left\lfloor\tfrac{1}{2}\delta n\right\rfloor-\lfloor c_{0}n\rfloor\,\Big|\,X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}=m\Big)\geq 1-\tfrac{\eta}{4},\quad\forall\,n\geq 1.$ (4.7)
@@
$\bar{h}:=h^{*}-\left\lfloor\tfrac{1}{2}\delta n\right\rfloor-\lfloor c_{0}n\rfloor.$
Combining (4.6) and (4.7), we arrive at the following high-probability event:
$\begin{aligned} &E(n,\delta,m):=\Big\{T\leq\bar{h},\,X_{h^{*}-\lfloor\frac{1}{2}\delta n\rfloor}=m\Big\}=\Big\{X_{k}=m,\ \forall\,k\in[\,\bar{h},\bar{h}+\lfloor c_{0}n\rfloor]\Big\},\\ &{\mathbb{P}}_{n}\bigg(\bigcup_{m=1}^{M_{0}}E(n,\delta,m)\bigg)\geq 1-\tfrac{\eta}{2}.\end{aligned}$ (4.8)
We shall take the event $E(n,\delta,m)$ as the typical event on which our subsequent analysis will be conditioned. Define for $1\leq m\leq M_{0}$ the conditional probability measure
${\mathbb{Q}}^{m}_{n}(\cdot):={\mathbb{P}}_{n}\big(\,\cdot\,\big|\,E(n,\delta,m)\big).$
Structure of $\mathscr{T}^{n}$ on $E(n,\delta,m)$
The purpose of this subsection is to set up the structural ingredients — coalescent subtrees, trunk paths and vertices, which will be crucial for the forthcoming estimates. Recall the subtree notation $\mathscr{T}^{n}(v)=\mathscr{T}^{n}(v,h^{*})$ from (3.17), which denotes the subtree of $\mathscr{T}^{n}$ rooted at $v$ and truncated at height $h^{*}$.
@@
$\Lambda_{j}:\mathbf{T}_{j}\longrightarrow\{1,\ldots,|\mathbf{T}_{j}|\},\qquad\Lambda_{j}(\mathsf{T}_{i}):=i,$ (4.9)
@@
In other words, for each $j$ between $\bar{h}$ and $\bar{h}+\lfloor c_{0}n\rfloor$, there are exactly $m$ vertices in $\mathcal{X}_{j}$. These vertices belong to $m$ distinct ancestral lineages that remain separate at height $\bar{h}$. Let $\mathsf{T}_{1},\ldots,\mathsf{T}_{m}$ be the $m$ coalescent subtrees at height $\bar{h}$. For $\bar{h}\leq j\leq\bar{h}+\lfloor c_{0}n\rfloor$ and $1\leq i\leq m$, denote by $z^{j}_{i}\in V(\mathsf{T}_{i})$ the unique vertex of $\mathsf{T}_{i}$ at height $j$. The sequence
@@
is called the trunk path of $\mathsf{T}_{i}$, and its elements are referred to as the trunk vertices of $\mathsf{T}_{i}$. The remaining $q_{n}(j)-m$ non-trunk vertices at height $j$ are denoted by $y^{j}_{1},\ldots,y^{j}_{q_{n}(j)-m}$, where the subscripts are assigned arbitrary.
@@
Assume (H1)–(H2). For fixed $\delta,s,\eta,M_{0},c_{0}$, there exists $\gamma>0$ such that for all $1\leq m\leq M_{0}$ and sufficiently large $n$,
${\mathbb{Q}}^{m}_{n}\big(\mathcal{N}_{i}\geq\gamma n^{2},\ \forall 1\leq i\leq m\big)\geq 1-\tfrac{\eta}{2}.$ (4.10)
Proof of (4.5) assuming Theorem 4.8.
We observe that each $\delta$-coalescent vertex at height $s\lfloor\delta n\rfloor$ has at least one descendant in $\{z^{0}_{1},\dots,z^{0}_{m}\}$. Hence, for all large $n$ and $1\leq m\leq M_{0}$,
${\mathbb{Q}}^{m}_{n}\big(B(n,s,\gamma,\delta)\big)\geq{\mathbb{Q}}^{m}_{n}\big(\mathcal{N}_{i}\geq\gamma n^{2},\ \forall 1\leq i\leq m\big)\geq 1-\tfrac{\eta}{2}.$
Combining with (4.8) completes the proof of (4.5). ∎
@@ 4.4.1 Outline of the proof
The first step is to partition the interval $\big[\bar{h},\bar{h}+\lfloor c_{0}n\rfloor\big]$ into $K_{\theta}$ sub-intervals $(h_{r},h_{r-1}]$, where $\bar{h}+\lfloor c_{0}n\rfloor=h_{0}>h_{1}>\cdots>h_{K_{\theta}}$ and each sub-interval has length $\lfloor\theta n\rfloor$ (see (4.11) below for precise definitions). For each $r=0,\ldots,K_{\theta}-1$, let $\mathcal{N}_{i,r}$ be the number of $y^{j}_{l}$ in the height interval $(h_{r},h_{r-1}]$ that merge into $Z_{i}$ before height $h_{r}$.
We then establish the independence of the family $(\mathcal{N}_{i,r})$ under $\mathbb{Q}^{m}_{n}$ (Proposition 4.9) and obtain a uniform lower bound for their tail probabilities (Proposition 4.10). Since the event that $\mathcal{N}_{i}>\gamma n^{2}$ is implied by the event that some $\mathcal{N}_{i,r}>\gamma n^{2}$, the independence and uniform lower bound together yield the desired probability estimate in (4.10).
@@
For these events, by using the general lemma (Lemma 4.12) for uniform selection pairs, we show that their probabilities can be decomposed into contributions from the genealogies above and below a given height. This decomposition then yields the desired independence property and simplifies the probability estimates required for proving the uniform lower bound.